A coin is thrown twice. The sample space is
S = (HH, HT, TH, TT,)
A = (HH, HT), B = (HT, TT)
Find the probability of A, B, AB and A+B.

Answer by robertb(5830)

P(A) = 1/2
P(B) = 1/2
P(AB) = P({HT}) = 1/4
P(A+B) = p({HH, HT, TT}) = 3/4